OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (anonymous):

Hi! I can help you! :)

OpenStudy (anonymous):

Do you have any ideas about how to get started, or are you completely stuck? :)

OpenStudy (johnweldon1993):

Hint...dot product \[\large \vec{u} \cdot \vec{v} = ||u|||v|| \cos(\theta)\] So \[\large cos(\theta) = \frac{\vec{u} \cdot \vec{v}}{||u|| ||v||}\]

OpenStudy (anonymous):

@BasketWeave completely stuck :-(

OpenStudy (anonymous):

Okay, no problem! John's given you a clue here: we can use the dot product.

OpenStudy (anonymous):

Do you know how to find the dot product between two vectors? :)

OpenStudy (anonymous):

u * v = <-5,-4> , <-4,-3> =-5(-4) + -4(-3) = 20 + 12 = 32 |u| = √(-5)^2 + (-4)^2 =√25 +16 = √41

OpenStudy (anonymous):

|v| = √(-4)^2 + (-3)^2 = √16 + 9 = √25 = 5

OpenStudy (anonymous):

u * v = 32 |u| =√41 |v| = 5

OpenStudy (anonymous):

\[\frac{ 32 }{ \sqrt{41}5} \]

OpenStudy (anonymous):

Well done :) Carry on...

OpenStudy (anonymous):

im getting 0.02648....

OpenStudy (anonymous):

i calculated wrong

OpenStudy (anonymous):

now i got, 0.99951.. one of my answer choices is 0.9 degrees, is that correct?

OpenStudy (anonymous):

Well, try and remember what that 0.99951 means... :) It's not actually the angle!

OpenStudy (johnweldon1993):

Right, that leaves us with \[\large cos(\theta) = 0.99951\] What do you get when you actually solve for the angle?

OpenStudy (anonymous):

1.8 degrees?

OpenStudy (anonymous):

Excellent, well done!

OpenStudy (johnweldon1993):

Correct indeed

OpenStudy (anonymous):

thank you guys so much for your help!

OpenStudy (anonymous):

You're most welcome :)